OpenStudy (anonymous):

A is 10% more than B, and C is 10% lesser than B. If the difference between A and C is 18, what is the value of A?

OpenStudy (anonymous):

a=110,b=100,c=90

OpenStudy (anonymous):

a-c=18

OpenStudy (anonymous):

110-90=20

OpenStudy (anonymous):

@dumbcow

OpenStudy (dumbcow):

you are on right track, just dont assume b=100 a = 1.1b c = 0.9b a-c = 18 1.1b - 0.9b = 18 .2b = 18 b = 90 --> a = 1.1*90